OpenStudy (anonymous):

What is the equation of the axis of symmetry for the function shown below? y-2=-(x+5)^2 A. y=2 B. X=5 c. x=2 d. x=-5

OpenStudy (anonymous):

axis of symmetry is for the x value. so you can rule out option A.

OpenStudy (anonymous):

formula is x = -b/2a

OpenStudy (anonymous):

convert it into a quadratic, or graph it to find the axis of symmetry.

OpenStudy (anonymous):

so the answer is 5?

OpenStudy (anonymous):

https://www20.csueastbay.edu/library/scaa/files/pdf/Conics.pdf Please take a look at Parabolas section, left side. y= a(x-h)^2 +k In your problem (x-h) = (x+5) x-h = x+5 h=-5 I think -5 is your answer. Hope the formula sheet will help in the future.

OpenStudy (anonymous):

thanks for the help :)

OpenStudy (anonymous):

Anytime :)